The probability of a new drug working in a patient in a clinical trial is modeled as P ∼ Unif([0, 1]). The drug is then given to patients one at a time, and you observe that the seventh patient is the first patient in which the drug works. That is, if X is a random variable denoting the first patient in which the drug worked, then [X|P] ∼ Geom(P), and you observed X = 7. Conditioned on this information that X = 7, what is the distribution of P?
Answer #1

el 8- Ocp4 Here xfp r Geometric (P). so, p(x=r|p) = (1-P) 12,221,2,3.0 Pwulo, o fep) = 1; O CPC = 0; elsewher so Joint DensitHere . g (plx=r) (1-3) . - B (2,8) Þ r Gengar from. There given that rat. gp|x=7) = + (1-) B/2,7) B/2.70 - 1(2) 17) - 1 6! os